The distance from the point P(1,3,2) to the line through (1,0,0), and (1,2,0) is given by ?


1
Expert's answer
2020-09-30T19:56:10-0400

Let M1(1, 0, 0) and M2(1, 2, 0).

We should find

"d=\frac{|\overrightarrow{PM_1}×\overrightarrow{s}|}{|\overrightarrow{s}|}" , where "\overrightarrow{s}=\overrightarrow{M_1M_2}=\{ 1-1, 2-0, 0-0\}=\{ 0, 2, 0\}" and "\overrightarrow{PM_1}=\{ 1-1, 0-3, 0-2\}=\{ 0, -3, -2\}" .

"\overrightarrow{PM_1}×\overrightarrow{s}=\begin{vmatrix}
 \overrightarrow{i} & \overrightarrow{j}&\overrightarrow{k} \\
 0 & -3&-2\\
0&2&0
\end{vmatrix}=\overrightarrow{i}(0+4)-\overrightarrow{j}(0-0)+\overrightarrow{k}(0-0)=4\overrightarrow{i}=\{ 4, 0, 0\}" .

Then "d=\frac{\sqrt{4^2+0^2+0^2}}{\sqrt{0^2+2^2+0^2}}=\frac{4}{2}=2" .